OpenStudy (anonymous):

I have an answer can someone just see if I'm right Rewrite the following function in the form f(x)=kx^p f(x)=(1/2x^3) k= p=

OpenStudy (anonymous):

k=.5 and p=-3

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

\[f(x)=kx^p\] \[p = -3\] \[k = 0.5\]

OpenStudy (anonymous):

so i am correct thanks

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

Did you have anymore questions?

OpenStudy (anonymous):

yes actually f(x)=3 square root x and then the 4 foes inside the check mark on the square root k= p=

OpenStudy (anonymous):

\[f(x)=3\sqrt[4]{x?}\]

OpenStudy (anonymous):

Ah yes, do you know what \[\sqrt[4]{x}\] is equivalent to?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sqrt{x} = x^{\frac{1}{2}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What would \[\sqrt[4]{x}\] be?

OpenStudy (anonymous):

so then does k= 3 and p= .25

OpenStudy (anonymous):

12.649

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

so k=3 and p=.25 right just making sure of the numbers

OpenStudy (anonymous):

yes, that's correct

OpenStudy (anonymous):

ok I have another so you want to check my answers for that too

OpenStudy (anonymous):

f(x)=(2/square root of x k= p=

OpenStudy (anonymous):

okay \[f(x)=\frac{2}{\sqrt{x}}\]

OpenStudy (anonymous):

yes k= 2 and p=1

OpenStudy (anonymous):

Do you know what happens when you bring the numerator to the denominator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sqrt{x}=x^{\frac{1}{2}}=\frac{1}{x^{-\frac{1}{2}}}\]

OpenStudy (anonymous):

so then k=2.5 and p=0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

k=2 and p=-0.5

OpenStudy (anonymous):

\[f(x)=\frac{2}{\sqrt{x}}=2*x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok thanks what about f(x)=(2x)(5 square eoot of x)

OpenStudy (anonymous):

k=10 and p=-1.5

OpenStudy (anonymous):

This? \[(2x)\sqrt[5]{x}\]

OpenStudy (anonymous):

no the 5 is next to the square root not in it

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

\[(2x)(5)\sqrt{x}\]

OpenStudy (anonymous):

\[(2x)(5\sqrt{x}\]

OpenStudy (anonymous):

This one is easy, we've already done something similar before

OpenStudy (anonymous):

k=10 and p=-1.5

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

\[2x(5*\sqrt{x})\] is equivalent to \[2x(5^1*x^{\frac{1}{2}})\]

OpenStudy (anonymous):

so then k=10 and p=1.5

OpenStudy (anonymous):

sorr yp=.5

OpenStudy (anonymous):

yes k=10 p=.5

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No problem.